Where did the cos(a) value came from?


1
Expert's answer
2021-01-16T17:21:10-0500

By the definition, in the right triangle the "cos\alpha" is the length of the adjacent side divided by the length of the hypotenuse:


"cos\alpha=\dfrac{Adjacent\ side}{Hypotenuse}."